OpenStudy (anonymous):

Evaluate 48 ÷ (6 + 2) • 3

OpenStudy (anonymous):

First do whats in the parenthesis

OpenStudy (anonymous):

6+2=8 so now are equation looks a bit like this 48/8*3

OpenStudy (anonymous):

BUT! Dont forget the order of operations which is PEMDAS orPEDMAS

OpenStudy (anonymous):

So in this case we do the problem straight across 48/8= what

OpenStudy (anonymous):

sammy

OpenStudy (anonymous):

do PEMDAS 48 ÷ (6 + 2) • 3 48 ÷ 8 • 3 6 • 3 = 18 18 is the answer

OpenStudy (anonymous):

Ugh

OpenStudy (anonymous):

I was geting to that i tryin to teach them XD

OpenStudy (anonymous):

@ParthKohli you did mistake 48/ (6 + 2) * 3 6+2 is 8 not subtract

OpenStudy (anonymous):

I said 8